Circular No. 6300

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1996 B2

Additional precise CCD observations:

1996 UT  α_{2000}  δ_{2000}  Observer
Feb. 1.11706 14^{h}32^{m}21.^{s}41 −24^{o}56′09′′.2 Tichý
1.12694 14 32 21.90 −24 56 08.9  _{”}
1.12980 14 32 22.05 −24 56 09.2  _{”}
1.21501 14 32 26.06 −24 56 11.5 Šarounová
1.21716 14 32 26.14 −24 56 11.2  _{”}
1.22469 14 32 26.47 −24 56 11.4  _{”}

M. Tichý (Kleť). 0.57-m f/5.2 reflector. L. Šarounová, P. Pravec, and M. Wolf (Ondřejov). 0.65-m reflector.

GRO J1744–28 W. Zhang, Goddard Space Flight Center (GSFC) and Universities

Space Research Association; E. Morgan, Massachusetts Institute of Technology; J. Swank and K. Jahoda, GSFC; and G. Jernigan and R. Klein, University of California at Berkeley, report: “GRO J1744–28 showed quasiperiodic intensity oscillations (QPO) in the energy band 3–12 keV during an 11-hr XTE observation starting on Jan. 18.77 UT. The power spectrum in the frequency band 5–100 Hz consists of a red-noise component that can be characterized by a power law with an index of −1 and three QPO peaks centered, respectively, at 20, 40, and 60 Hz. The peaks at 20 and 60 Hz are clearly discernible, yet hard to quantify at this stage of the analysis. The peak at 40 Hz can be well characterized by a Lorentzian profile with a FWHM of 27 Hz. The total integrated power under the 40-Hz peak amounts to a fractional RMS variation of 5.6 percent (not corrected for the effect of detector deadtime). The effect of flares on the centroid frequencies and the strength of the oscillations is currently under study. It is not clear whether the standard beat frequency model can accommodate this QPO because of the high frequencies. The pulsation at 2 Hz (IAUC 6285) and the energy spectrum (IAUC 6291) imply a strong magnetic field, which probably prevents the accretion disk from extending so close to the neutron star as to have keplerian frequencies as high as the observed QPO frequencies, even if one takes into account the presumably very high luminosity.”

1996 February 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green